OpenStudy (anonymous):

Write the expression for the phrase. the quantity two times d minus five times the quantity six times d plus 4. I know this is 2d - 5 and 6d + 4 But where do I go from here..? Is it (2d - 5) * (6d + 4) ? @shamil98 @SolomonZelman

OpenStudy (anonymous):

Like can't I combine them?

OpenStudy (solomonzelman):

\[2-d-5(6d+4)\]

OpenStudy (solomonzelman):

yes, expand the parenthesis and add like terms.

OpenStudy (anonymous):

Wait where did you get 2 - d?

OpenStudy (solomonzelman):

I missed the words, "the quantity", sorry.

OpenStudy (solomonzelman):

\[2(d-5) \times 6(d+4)\]

OpenStudy (solomonzelman):

\[(2d-10) \times (6d+24)\]\[OR:~~~~~~~~~(2d-10)(6d+24)\]EXPAND:

OpenStudy (solomonzelman):

\[OR:~~~12 \times (d-5)(d+4)\]same thing, which ever one you like....

OpenStudy (anonymous):

Oh I see. And it's okay, we all make mistakes, haha.

OpenStudy (solomonzelman):

Alright....